OpenStudy (anonymous):

square root of 50 minus square root of 18

OpenStudy (joftheworld):

2.82

OpenStudy (praxer):

$$\sqrt50-\sqrt18 \\= 5\sqrt2-3\sqrt2\\=2\sqrt2$$ $$\explanation$$ : In this type of question try to factorise the number under the roots lets take for eg: 50=5*5*2 which is $$\sqrt{5^2*2}\\=\sqrt{5^2}.\sqrt2\\\ we \ know\ that\sqrt{5^2} \ is \ 5. $$ hence do the same for the next and you will find the solution. I hope I was able to give the explanation well.

OpenStudy (praxer):

@Andrewlanni Do check the solution.